A triode valve operates at V^p = 225 V and V^g = -0.5 V. The plate current remains unchanged if the plate voltage is increased to 250 V and the grid voltage is decreased to -2.5 V. Calculate the amplification factor.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Amplification factor for triode value = μ = Charge in Plate Voltage/Charge in Grid Voltage = δV base p/ δV base g = 250 – 225/2.5 – 0.5 = 25/2 = 12.5 [∴ δV base p = 250 – 225, δ base g = 2.5 – 0.5]